# CarbonClarity: Understanding and Addressing Uncertainty in Embodied Carbon for Sustainable Computing

### v1

@@ Abstract
Index Terms:
sustainable computing, manufacturing, chiplet, embodied carbon, uncertaintyI Introduction
@@
Upon publication, we intend to open-source CarbonClarity to enable robust, sustainable hardware design.
@@ IV-A System Provisioning Using TPUs as An Example
Fig. 8 (right) also illustrates that while reducing the array size decreases embodied carbon and its uncertainty, it leads to diminishing performance returns. Companies optimizing purely for worst-case carbon compliance would prefer the 32×32 design, as it minimizes risk but comes at a major performance cost. In contrast, using the full uncertainty distribution enables companies to adopt a more balanced strategy: selecting a 128×128 TPU size improves performance by 34% compared to the conservative 32×32 option, while keeping the probability of exceeding the carbon target below 5%. Conversely, relying on average-case estimates alone and choosing the 256×256 design would risk a 21.2% carbon overrun. Thus, uncertainty-aware modeling helps companies provision hardware aligned with both performance and sustainability goals..
@@ V-C Case Study III: Identifying Key Sources of Uncertainty
Yield uncertainty creates a long-tail effect for large-area chips, as shown in Fig. 13c. Larger chips not only generate more embodied carbon but also exhibit greater uncertainty in their emissions. The long-tail effect of yield helps explain why the chiplet design in Case Study I effectively reduces its carbon emissions, especially compared to monolithic chips with large area.

### v2

@@ Abstract
I Introduction
@@
CarbonClarity is available in the below GitHub repository: https://github.com/S4AI-CornellTech/CarbonClarity
@@ IV-A System Provisioning Using TPUs as An Example
Fig. 8 (right) also illustrates that while reducing the array size decreases embodied carbon and its uncertainty, it leads to diminishing performance returns. Companies optimizing purely for worst-case carbon compliance would prefer the 64×64 design, as it minimizes risk but comes at a major performance cost. In contrast, using the full uncertainty distribution enables companies to adopt a more balanced strategy: selecting a 128×128 TPU size improves performance by 34% compared to the conservative 64×64 option, while keeping the probability of exceeding the carbon target below 5%. Conversely, relying on average-case estimates alone and choosing the 256×256 design would risk a 21.2% carbon overrun. Thus, uncertainty-aware modeling helps companies provision hardware aligned with both performance and sustainability goals..
@@ V-C Case Study III: Identifying Key Sources of Uncertainty
Yield uncertainty creates a long-tail effect for large-area chips, as shown in Fig. 13c. Larger chips not only generate more embodied carbon but also exhibit greater uncertainty in their emissions. The long-tail effect of yield helps explain why the chiplet design in Case Study I effectively reduces its carbon emissions, especially compared to large monolithic chips.
@@ VI Conclusion and Future Works
VII Acknowledgment
This research was supported by the National Science Foundation under award numbers CCF-2324860 and CNS-2335795.